OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using substitution: -x + 6y − 10 = 0 3x − 6y = 1 Replace the right side of the second equation with the first equation. Write the first equation in terms of x. Multiply the second equation by 3. Multiply the first equation by 3.

OpenStudy (anonymous):

help @AlwaystheBookworm

OpenStudy (anonymous):

Multiply the first equation by 3

OpenStudy (anonymous):

Oh! Wait, it said substitution. I answered according to elimination

OpenStudy (anonymous):

i need to know if this is the correct answer for this question Solve using elimination: -4x = 2 + 2y -7x − 2y = -4 (2, -5) Infinitely many solutions No solution (3, -7) is (3,-7) the correct answer

OpenStudy (anonymous):

That is correct

OpenStudy (anonymous):

o ok :/ so would it be multiply the second equation by 3 and ok

OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using substitution: 7x + -y − 5 = 0 3x − 4y = 2 Write the first equation in terms of y. Multiply the second equation by -7. Replace the right side of the second equation with the first equation. Multiply the first equation by -4. .is multipliying the first equation by -4 the right answer

OpenStudy (anonymous):

No, because that would be eliminating the y. That's using the process of elimination. You want substitution.

OpenStudy (anonymous):

replace the right side of the second equation with the first equation is the correct answer

OpenStudy (anonymous):

That's correct. I'm not so great at explaining the substitution method, but here's a site that does it pretty well. http://www.purplemath.com/modules/systlin4.htm

OpenStudy (anonymous):

thanks i have 3 more question you dont mind helping me do you

OpenStudy (anonymous):

Solve using substitution: -x + 4y = -9 12y − 3x = 36 Infinitely many solutions No solution (-1, -2) (5,-1) is (5,-1) the right answer